OpenStudy (anonymous):

1. Solve the formula for the area of a triangle A = (1/2) bh for height h.

OpenStudy (anonymous):

what are the number

OpenStudy (anonymous):

no mtter for numbers

OpenStudy (anonymous):

to solve for h divide A by ((1/2)b) h=a/1/2/b=2a/b

OpenStudy (anonymous):

yea im in the test now